# How Global Calibration Strengthens Multiaccuracy

### v1

@@ 1 Introduction
Multicalibrated predictors with respect to a family $\mathcal{C}$ can be learned from a sample of individuals and outcome pairs via a learning primitive known as weak agnostic learning for $\mathcal{C}$ [24]. Weak agnostic learning allows learning a hypothesis that is somewhat correlated with the labels as long as some hypothesis in $\mathcal{C}$ is sufficiently correlated [7, 34, 30, 13]. Quantitatively, assuming the existence of a hypothesis in $\mathcal{C}$ that achieves correlation $\alpha>0$, an $(\alpha,\beta)$ weak learner returns a hypothesis with correlation at least $\beta$. Typically, one hopes to get $\beta=\operatorname{poly}(\alpha)$ even when $\alpha$ approaches 00. This should be contrasted with (strong) agnostic learning, where a hypothesis can be found that is almost as correlated with the data as the best hypothesis in the class $\mathcal{C}$ is; in other words $\beta\geq\alpha-\varepsilon$.
@@
  • Perhaps surprisingly, we give evidence that multiaccuracy in itself need not imply weak agnostic learning in the standard sense. We prove the existence of a class $\mathcal{C}$, a distribution on labels, and a $\mathcal{C}$-multiaccurate predictor $p$ such that even though there is hypothesis in $\mathcal{C}$ that gives correlation $1/2$, no post-processing of the predictor has any non-zero correlation with the labels.
  • We complement this lower bound by showing that multiaccuracy does imply weak learning once the correlation $\alpha$ exceeds $1/2$: If there is a hypothesis in $\mathcal{C}$ with correlation $\alpha$, then $\mathrm{sign}(p-1/2)$ where $p$ is $(\mathcal{C},0)$-multiaccurate achieves correlation $\beta=2\alpha-1$, a guarantee that is only non-trivial when $\alpha>1/2$. We call this restricted weak agnostic learning, to contrast it with the standard notion of weak agnostic learning where $\alpha$ approaches 00.
@@
  • Lastly, the result of [24] shows that $(\alpha,\beta)$-weak agnostic learning achieves a multiaccuracy parameter of $\alpha$; we show that in general, given an such a weak learner, it is not possible to achieve multiaccuracy parameter better than $\alpha/2$.
@@
Impagliazzo’s hardcore lemma is a fundamental result in computational complexity that relates a function $g$ being (mildly) $\delta$-hard to predict under a distribution $\mathcal{D}$ for distinguishers in a class $\mathcal{C}_{q}$ to the existence of a hardcore measure $\mu$ of density $\Omega(\delta)$ in $\mathcal{D}$, such that $g$ is $1/2-\varepsilon$-hard to predict (i.e., cannot improve over random guessing) when sampling from the measure to distinguishers from a family $\mathcal{C}\subset\mathcal{C}_{q}$. The parameter $q$ measures the number of $\mathcal{C}$-gates required for functions in $\mathcal{C}_{q}$. The larger $q$ is, the stronger the hardness assumption we make to show the existence of a hardcore distribution for $\mathcal{C}$.
The work of Trevisan, Tulsiani, and Vadhan [46], which preceded the literature on multigroup fairness, gave an elegant connection to multiaccuracy, showing how to get a simple and explicit hardcore measure from a $\mathcal{C}$-multiaccurate predictor. [^2] As pointed out in [11, 8], the notion of $(\mathcal{C},\epsilon)$-indistinguishability in complexity theory, as used in [46], is exactly equivalent to the notion of $(\mathcal{C},\epsilon)$-multiaccuracy as defined in the multigroup fairness literature. Their construction, however, achieves a hardcore measure of density $\delta$, while the optimal density is $2\delta$ [26]. Recently, Casacuberta, Dwork, and Vadhan [8] showed that multicalibration implies a stronger form of the hardcore lemma, which they call ICHL+++++ +. From this they derive the original hardcore lemma by Impagliazzo, this time with a hardcore measure of optimal density $2\delta$. However, while [46] achieve $q=O(1/\varepsilon^{2}\delta^{2})$, [8] get $q=O(1/(\varepsilon^{12}\delta^{6}))$. [^3] The stronger bound of $O(1/(\varepsilon^{2}\delta^{2}))$ stated in [8] appears not to be correct. We refer the reader to a detailed discussion at the end of Section 5.3. This increased complexity comes in part because achieving multicalibration requires more calls to a (strong) agnostic learner for $\mathcal{C}$, in contrast to multiaccuracy.
@@
  • Multicalibration: the top tier. Multicalibration implies the strongest known form of the Impagliazzo hardcore lemma, namely the IHCL+++++ + [8]. On the learning front, it implies several strong forms of learning: swap agnostic learning and swap omniprediction (we refer the readers to [17, 18] for the definitions of these notions). It however comes at the cost of higher complexity of learning, more complex predictors (in terms of circuit size) and higher sample complexity.
  • Calibrated Multiaccuracy: the Goldilocks zone. On the learning front, calibrated multiaccuracy implies strong agnostic learning and loss OI for generalized linear models (see [16] for definitions). On the hardcore lemma front, it gives hardcore sets of optimal density $2\delta$, but with relatively low blowup in $q$. The computational, and representational, costs are much lower than multicalibration, and comparable to multiaccuracy. But it is not as powerful as full-blown multicalibration, and does not imply IHCL+++++ + or swap agnostic learning (see [18] for the latter separation).
@@
The power of multicalibration [24, 10, 17] illustrates the potency of combining calibration and multiaccuracy. But this great power comes at a high computational price. Our results show that in two important settings, agnostic learning and hardcore set constructions, global calibration together with multiaccuracy is a cheaper alternative to multicalibration, adding weight to a thesis that was first put forth by [16]. Our proofs highlight the complementary roles that multiaccuracy and calibration play, and give clear intuition for why their combination succeeds even when they each fail in isolation. A deep insight in theoretical computer science has been that hardness and learning are two sides of the same coin [27, 35, 46, 13]; our work provides further evidence of this phenomenon.
@@ 2 Overview of our Results
Calibrated multiaccuracy introduced in [16], asks for predictors that are both multiaccurate and calibrated, as the name suggests. It is a stronger requirement than just multiaccuracy, although of comparable sample and computational complexity [16, Section 7].
@@
$\max_{c\in\mathcal{C}}\operatorname*{\mathbf{E}}[|c(\mathbf{x})(\operatorname*{\mathbf{E}}[\mathbf{y}|p(\mathbf{x})]-p(\mathbf{x}))|]\leq\tau.$
@@
We define weak and strong agnostic learning following [7, 36, 34, 30, 13]. The correlation between a function $c:\mathcal{X}\to[-1,1]$ and the labels is defined as $\mathrm{cor}_{\mathcal{D}}(\mathbf{y},c)=\operatorname*{\mathbf{E}}[c(\mathbf{x})(2\mathbf{y}-1)]\in[-1,1]$, where the map $\mathbf{y}\mapsto 2\mathbf{y}-1$ maps labels in $\{0,1\}$ to {±$\{\pm 1\}$. A $\beta$-weak learner is a function $h:\mathcal{X}\to[-1,1]$ such that $\mathrm{cor}(\mathbf{y},h)\geq\beta$. Taking $h=0$ which gives 00 correlation as our baseline, we would like $\beta\in(0,1]$ to be as large as possible. For a predictor $p:\mathcal{X}\to\{0,1\}$, we will often use the transformation $p\mapsto 2p-1$. This maps the $p=1/2$ predictor to $h=0$, corresponding to the fact that guessing randomly at every point gives 00 correlation.
For $\alpha\geq\beta\in[0,1]$, an $(\alpha,\beta)$ weak agnostic learner for $\mathcal{C}$ is an algorithm that has access to a distribution $\mathcal{D}$ (with a known marginal $\mathcal{D}_{\mathcal{X}}$ over $\mathcal{X}$). If $\mathcal{D}$ is such that if there exists a hypothesis $c$ such that $\mathrm{cor}(\mathbf{y},c)\geq\alpha$, the algorithm is guaranteed to return a $\beta$-weak learner. It helps to think of $\beta$ as smaller than but polynomially related to $\alpha$. In contrast, a strong agnostic learner must return a hypothesis that achieves correlation $\mathrm{Opt}(\mathcal{C})-\varepsilon$ where $\mathrm{Opt}(\mathcal{C})=\max_{c\in\mathcal{C}}\mathrm{cor}(\mathbf{y},c)$ for every $\varepsilon>0$. The agnostic boosting results of [30, 13] imply that an $(\alpha,\beta)$ weak learner for $\mathcal{C}$ can be boosted to a hypothesis that gives correlation at least $\mathrm{Opt}(\mathcal{C})-\alpha-\varepsilon$. Thus, given a weak learner that can handle arbitrarily small correlation $\alpha$, we can achieve strong agnostic learning. [^5] The parameter $\beta$ only affects the number of iterations of boosting, hence it is less important than $\alpha$.
@@
Results of [24, 16] show that all of multiaccuracy, calibrated multiaccuracy, and multicalibration with error $\tau$ can be achieved using an $(\tau,\beta)$ weak agnostic learner as an oracle. The algorithms will involve multiple calls to the weak agnostic learner. The number of calls grows as $O(1/\beta^{2})$ for multiaccuracy and calibrated multiaccuracy, and $O(1/\tau^{2}\beta^{4})$ for multicalibration. The added complexity in multicalibration comes from having to ensure multiaccuracy holds for prediction values that have fairly small probability, see the detailed discussion in [16, Section 7.3].
@@
  • When $\mathcal{C}=\{c:\mathcal{X}\to\{\pm 1\}\}$ is a family of Boolean rather than bounded functions, then correlation coincides with the 00-$1$ loss. In this setting, the work of [16] shows that if $p$ satisfies calibrated multiaccuracy, then it is an omnipredictor, using the technique of loss outcome indistinguishability. This implies that the hypothesis $\mathrm{sign}(p-1/2)$ is a strong agnostic learner.
@@ 2.1 Agnostic Learning and Multiaccuracy: With and Without Calibration
then $12h(x)-$ 1 is a $2\gamma$ weak learner. A similar result is proved in the work of [2]. With this characterization in hand, we can rephrase our question to ask when a multiaccurate predictor improves upon random guessing in terms of squared loss. More generally, for what values of $\alpha,\beta$ can one get an $(\alpha,\beta)$-weak learner for $\mathcal{C}$ by post-processing a $\mathcal{C}$-multiaccurate predictor?
@@
Motivated by Theorem 3.5, let us first focus on the hypothesis $q(x)=2p(x)-1$ and understand when this fails as a weak learner. To frame this as a geometric question, let $q^{*}(x)=2p^{*}(x)-1$. Assume there exists $c\in\mathcal{C}$ so that $\langle q^{*},c\rangle=\mathrm{cor}(\mathbf{y},c)\geq\alpha$. By multiaccuracy, this also means that $\langle q,c\rangle\geq\alpha$, since multiaccuracy preserves the projection onto $\mathcal{C}$. For what values of $\alpha$ is it possible that $\langle q^{*},q\rangle=0$, so that $q$ is not a $\beta$-weak learner for $\mathcal{C}$?
We will show that the answer is $\alpha\leq 1/2$. Take $q,q^{*}$ to be vectors in {±$\{\pm 1\}^{N}$, so that $\langle q,q^{*}\rangle=0$. Let $d_{H}$ denote the normalized Hamming distance in this space. Let $c\in\{\pm 1\}^{n}$ be the mid-point of $q,q^{*}$ in the Hamming metric. Since $d_{H}(q,c)=d_{H}(q^{*},c)=1/4$, we have $\langle q^{*},c\rangle=\langle q,c\rangle=1/2$. Thus we have a triple of vectors $q^{*},q,c\in\{\pm 1\}^{N}$ such that $\langle q^{*},c\rangle=\langle q,c\rangle=1/2$, whereas $\langle q^{*},q\rangle=0$. Translating this geometric example to a learning setting requires some care, but it conveys the intuition for why $1/2$ arises as the right threshold.
@@
The key challenge is ruling out all post-processing functions $k(p)$, not just $12p-$ 1. Our construction achieves this by ensuring that for every predicted value $v$, $\operatorname*{\mathbf{E}}[\mathbf{y}|p(\mathbf{x})=v]=1/2$. Thus the predictor is anti-calibrated in the sense that even conditioned on its predictions, the bit $\mathbf{y}$ is uniformly random. This property ensures that no function $k(p)$ can beat random guessing.
@@
Complementing this lower bound, Theorem 4.2 shows that $p$ being $\mathcal{C}$-multiaccurate implies that $12p-$ 1 is an $(\alpha,2\alpha-1)$ weak agnostic learner. Note that $\beta>0$ only when $\alpha>1/2$. This is in contrast with the standard usage of weak learning, where on typically thinks of $\alpha$ being close to 00, hence we refer to this notion as restricted weak agnostic learning. In the realizable setting where $\alpha=1$, we get $\beta=1$.
Intuition for why this is the right bound on $\beta$ can again be obtained from the geometric setting, where $q^{*},q,c$ are vectors in {±$\{\pm 1\}^{N}$. The condition $\langle q,c\rangle\geq\alpha$ translates to $d_{H}(q,c)\leq(1-\alpha)/2$, and similarly for $d_{H}(q^{*},c)$. By the triangle inequality, we must have
@@
While translating this intuition from the Boolean to the continuous setting is by no means immediate, we show that $12p-$ 1 is indeed a $\beta=2\alpha-1$-weak learner. The key step is to show the following upper bound on the correlation obtainable using any hypothesis in $\mathcal{C}$: if $p$ is $(\mathcal{C},\tau)$-multiaccurate, then
@@
Thus a calibrated predictor that deviates noticeably from random guessing (saying $1/2$ all the time) gives rise to an agnostic learner. But if we add the assumption that $p$ is $\mathcal{C}$-multiaccurate, and that there exists $c\in\mathcal{C}$ so that $\mathrm{cor}(\mathbf{y},c\mathbf{x})\geq\alpha$, then $p$ can no longer predict $1/2$ all the time. This is proved in Lemma 4.4 which shows that for every $c\in\mathcal{C}$,
@@
We note that this result can also be shown using the loss OI technique of [16], where it is applied to show agnostic learning for Boolean function classes $\mathcal{C}$ (where correlation coincides with 00-$1$ loss). But we find the proof above more intuitive as it highlights how multiaccuracy and calibration complement each other.

### v2

@@ 1 Introduction
Multicalibrated predictors with respect to a family $\mathcal{C}$ can be learned from a sample of individuals and outcome pairs via a learning primitive known as weak agnostic learning for $\mathcal{C}$ [24]. Weak agnostic learning allows learning a hypothesis that is somewhat correlated with the labels as long as some hypothesis in $\mathcal{C}$ is sufficiently correlated [3, 31, 30, 13]. Quantitatively, assuming the existence of a hypothesis in $\mathcal{C}$ that achieves correlation $\alpha>0$, an $(\alpha,\beta)$ weak learner returns a hypothesis with correlation at least $\beta$. Typically, one hopes to get $\beta=\operatorname{poly}(\alpha)$ even when $\alpha$ approaches 0. This should be contrasted with (strong) agnostic learning, where a hypothesis can be found that is almost as correlated with the data as the best hypothesis in the class $\mathcal{C}$ is; in other words $\beta\geq\alpha-\varepsilon$.
@@
  • Perhaps surprisingly, we give evidence that multiaccuracy in itself need not imply weak agnostic learning in the standard sense. We prove the existence of a class $\mathcal{C}$, a distribution on labels, and a $\mathcal{C}$-multiaccurate predictor $p$ such that even though there is hypothesis in $\mathcal{C}$ that gives correlation $1/2$, no post-processing of the predictor has any non-zero correlation with the labels. Furthermore, we show that this is the case for all classes $\mathcal{C}$ for which there exist of pseudorandom functions that fool $\mathcal{C}$.
  • We complement this lower bound by showing that multiaccuracy does imply weak learning once the correlation $\alpha$ exceeds $1/2$: If there is a hypothesis in $\mathcal{C}$ with correlation $\alpha$, then $\mathrm{sign}(p-1/2)$ where $p$ is $(\mathcal{C},0)$-multiaccurate achieves correlation $\beta=2\alpha-1$, a guarantee that is only non-trivial when $\alpha>1/2$. We call this restricted weak agnostic learning, to contrast it with the standard notion of weak agnostic learning where $\alpha$ approaches 0.
@@
  • Lastly, the result of [24] shows that $(\alpha,\beta)$-weak agnostic learning achieves a multiaccuracy parameter of $\alpha$; we show that in general, given such a weak learner, it is not possible to achieve multiaccuracy parameter better than $\alpha/2$.
@@
Impagliazzo’s hardcore lemma is a fundamental result in computational complexity that relates a function $g$ being (mildly) $\delta$-hard to predict under a distribution $\mathcal{D}$ for distinguishers in a class $\mathcal{C}_{q}$ to the existence of a hardcore measure $\mu$ of density $\Omega(\delta)$ in $\mathcal{D}$, such that $g$ is $1/2-\varepsilon$-hard to predict (i.e., one cannot improve over random guessing) when sampling from the measure to distinguishers from a family $\mathcal{C}\subset\mathcal{C}_{q}$. The parameter $q$ measures the number of $\mathcal{C}$-gates required for functions in $\mathcal{C}_{q}$. The larger $q$ is, the stronger the hardness assumption we make to show the existence of a hardcore distribution for $\mathcal{C}$.
The work of Trevisan, Tulsiani, and Vadhan [46], which preceded the literature on multigroup fairness, gave an elegant connection to multiaccuracy, showing how to get a simple and explicit hardcore measure from a $\mathcal{C}$-multiaccurate predictor. [^2] As pointed out in [11, 8], the notion of $(\mathcal{C},\epsilon)$-indistinguishability in complexity theory, as used in [46], is exactly equivalent to the notion of $(\mathcal{C},\epsilon)$-multiaccuracy as defined in the multigroup fairness literature. Their construction, however, achieves a hardcore measure of density $\delta$, while the optimal density is $2\delta$ [25]. Recently, Casacuberta, Dwork, and Vadhan [8] showed that multicalibration implies a stronger form of the hardcore lemma, which they call ICHL++++. From this they derive the original hardcore lemma by Impagliazzo, this time with a hardcore measure of optimal density $2\delta$. However, while [46] achieve $q=O(1/\varepsilon^{2}\delta^{2})$, [8] get $q=O(1/(\varepsilon^{12}\delta^{6}))$. [^3] The stronger bound of $O(1/(\varepsilon^{2}\delta^{2}))$ stated in [8] appears not to be correct. We refer the reader to a detailed discussion at the end of Section 5.3. This increased complexity comes in part because achieving multicalibration requires more calls to a (strong) agnostic learner for $\mathcal{C}$, in contrast to multiaccuracy.
@@
  • Multicalibration: the top tier. Multicalibration implies the strongest known form of the Impagliazzo hardcore lemma, namely the IHCL++++ [8]. On the learning front, it implies several strong forms of learning: swap agnostic learning and swap omniprediction (we refer the readers to [18, 20] for the definitions of these notions). It however comes at the cost of higher complexity of learning, more complex predictors (in terms of circuit size) and higher sample complexity.
  • Calibrated Multiaccuracy: the Goldilocks zone. On the learning front, calibrated multiaccuracy implies strong agnostic learning and loss OI for generalized linear models (see [17] for definitions). On the hardcore lemma front, it gives hardcore sets of optimal density $2\delta$, but with relatively low blowup in $q$. The computational, and representational, costs are much lower than multicalibration, and comparable to multiaccuracy. But it is not as powerful as full-blown multicalibration, and does not imply IHCL++++ or swap agnostic learning (see [20] for the latter separation).
@@
The power of multicalibration [24, 10, 18] illustrates the potency of combining calibration and multiaccuracy. But this great power comes at a high computational price. Our results show that in two important settings, agnostic learning and hardcore set constructions, global calibration together with multiaccuracy is a cheaper alternative to multicalibration, adding weight to a thesis that was first put forth by [17]. Our proofs highlight the complementary roles that multiaccuracy and calibration play, and give clear intuition for why their combination succeeds even when they each fail in isolation. In concurrent work, Hu and Vadhan use the notion of calibrated multiaccuracy to provide pseudoentropy characterizations for general entropy notions [27]. A deep insight in theoretical computer science has been that hardness and learning are two sides of the same coin [28, 35, 46, 13]; our work provides further evidence of this phenomenon.
@@ 2 Overview of our Results
Calibrated multiaccuracy introduced in [17], asks for predictors that are both multiaccurate and calibrated, as the name suggests. It is a stronger requirement than just multiaccuracy, although of comparable sample and computational complexity [17, Section 7].
@@
[…]
@@
We define weak and strong agnostic learning following [3, 32, 31, 30, 13]. The correlation between a function $c:\mathcal{X}\to[-1,1]$ and the labels is defined as $\mathrm{cor}_{\mathcal{D}}(\mathbf{y},c)=\operatorname*{\mathbf{E}}[c(\mathbf{x})(2\mathbf{y}-1)]\in[-1,1]$, where the map $\mathbf{y}\mapsto 2\mathbf{y}-1$ maps labels in $\{0,1\}$ to $\{\pm 1\}$. A $\beta$-weak learner is a function $h:\mathcal{X}\to[-1,1]$ such that $\mathrm{cor}(\mathbf{y},h)\geq\beta$. Taking $h=0$ which gives 0 correlation as our baseline, we would like $\beta\in(0,1]$ to be as large as possible. For a predictor $p:\mathcal{X}\to\{0,1\}$, we will often use the transformation $p\mapsto 2p-1$. This maps the $p=1/2$ predictor to $h=0$, corresponding to the fact that guessing randomly at every point gives 0 correlation.
For $\alpha\geq\beta\in[0,1]$, an $(\alpha,\beta)$-weak agnostic learner for $\mathcal{C}$ is an algorithm that has access to a distribution $\mathcal{D}$ (with a known marginal $\mathcal{D}_{\mathcal{X}}$ over $\mathcal{X}$). If $\mathcal{D}$ is such that if there exists a hypothesis $c$ such that $\mathrm{cor}(\mathbf{y},c)\geq\alpha$, the algorithm is guaranteed to return a $\beta$-weak learner. It helps to think of $\beta$ as smaller than but polynomially related to $\alpha$. In contrast, a strong agnostic learner must return a hypothesis that achieves correlation $\mathrm{Opt}(\mathcal{C})-\varepsilon$ where $\mathrm{Opt}(\mathcal{C})=\max_{c\in\mathcal{C}}\mathrm{cor}(\mathbf{y},c)$ for every $\varepsilon>0$. The agnostic boosting results of [30, 13] imply that an $(\alpha,\beta)$ weak learner for $\mathcal{C}$ can be boosted to a hypothesis that gives correlation at least $\mathrm{Opt}(\mathcal{C})-\alpha-\varepsilon$. Thus, given a weak learner that can handle arbitrarily small correlation $\alpha$, we can achieve strong agnostic learning. [^5] The parameter $\beta$ only affects the number of iterations of boosting, hence it is less important than $\alpha$.
@@
Results of [24, 17] show that all of multiaccuracy, calibrated multiaccuracy, and multicalibration with error $\tau$ can be achieved using an $(\tau,\beta)$ weak agnostic learner as an oracle. The algorithms will involve multiple calls to the weak agnostic learner. The number of calls grows as $O(1/\beta^{2})$ for multiaccuracy and calibrated multiaccuracy, and $O(1/\tau^{2}\beta^{4})$ for multicalibration. The added complexity in multicalibration comes from having to ensure multiaccuracy holds for prediction values that have fairly small probability, see the detailed discussion in [17, Section 7.3].
@@
  • When $\mathcal{C}=\{c:\mathcal{X}\to\{\pm 1\}\}$ is a family of Boolean rather than bounded functions, then correlation coincides with the 0-11 loss. In this setting, the work of [17] shows that if $p$ satisfies calibrated multiaccuracy, then it is an omnipredictor, using the technique of loss outcome indistinguishability. This implies that the hypothesis $\mathrm{sign}(p-1/2)$ is a strong agnostic learner.
@@ 2.1 Agnostic Learning and Multiaccuracy: With and Without Calibration
then 2h(x)12h(x)-1 is a $2\gamma$ weak learner. A similar result is proved in the work of [7]. With this characterization in hand, we can rephrase our question to ask when a multiaccurate predictor improves upon random guessing in terms of squared loss. More generally, for what values of $\alpha,\beta$ can one get an $(\alpha,\beta)$-weak learner for $\mathcal{C}$ by post-processing a $\mathcal{C}$-multiaccurate predictor?
@@
Motivated by Theorem 3.5, let us first focus on the hypothesis q(x)=2p(x)1q(x)=2p(x)-1 and understand when this fails as a weak learner. To frame this as a geometric question, let $q^{*}(x)=2p^{*}(x)-1$. Assume there exists $c\in\mathcal{C}$ so that $\langle q^{*},c\rangle=\mathrm{cor}(\mathbf{y},c)\geq\alpha$. By multiaccuracy, this also means that $\langle q,c\rangle\geq\alpha$, since multiaccuracy preserves the projection onto $\mathcal{C}$. For what values of $\alpha$ is it possible that $\langle q^{*},q\rangle=0$, so that $q$ is not a $\beta$-weak learner for $\mathcal{C}$?
We will show that the answer is $\alpha\leq 1/2$. Take $q,q^{*}$ to be vectors in $\{\pm 1\}^{n}$, so that $\langle q,q^{*}\rangle=0$. Let $d_{H}$ denote the normalized Hamming distance in this space. Let $c\in\{\pm 1\}^{n}$ be the mid-point of $q,q^{*}$ in the Hamming metric. Since $d_{H}(q,c)=d_{H}(q^{*},c)=1/4$, we have $\langle q^{*},c\rangle=\langle q,c\rangle=1/2$. Thus we have a triple of vectors $q^{*},q,c\in\{\pm 1\}^{N}$ such that $\langle q^{*},c\rangle=\langle q,c\rangle=1/2$, whereas $\langle q^{*},q\rangle=0$. Translating this geometric example to a learning setting requires some care, but it conveys the intuition for why $1/2$ arises as the right threshold.
@@
The key challenge is ruling out all post-processing functions $k(p)$, not just 2p12p-1. Our construction achieves this by ensuring that for every predicted value $v$, $\operatorname*{\mathbf{E}}[\mathbf{y}|p(\mathbf{x})=v]=1/2$. Thus the predictor is anti-calibrated in the sense that even conditioned on its predictions, the bit $\mathbf{y}$ is uniformly random. This property ensures that no function $k(p)$ can beat random guessing.
Our construction in Theorem 4.1 makes careful use of the properties of majority and parity functions with respect to the uniform distribution. However, under the additional assumption of the existence of pseudorandom functions, we show in Theorem 4.2 that for any reasonable class $\mathcal{C}$, a multiaccurate predictor cannot in general be post-processed to have non-trivial correlation with the labels, even when $\mathrm{Opt}(\mathcal{C})$ is large. The key idea is based on making $p^{*}$ and $p$ both be averages of a function $f$ from $\mathcal{C}$ and a pseudorandom (for $\mathcal{C}$) function $g$, but with opposite signs. Thus, $p^{*}-p$ is essentially a pseudorandom function, which ensures multiaccuracy. However, even though $p^{*}$ has a high correlation with some $c\in\mathcal{C}$, $\mathrm{cor}(p^{*},p)\approx 0$.
@@
Complementing this lower bound, Theorem 4.4 shows that $p$ being $\mathcal{C}$-multiaccurate implies that 2p12p-1 is an $(\alpha,2\alpha-1)$ weak agnostic learner. Note that $\beta>0$ only when $\alpha>1/2$. This is in contrast with the standard usage of weak learning, where one typically thinks of $\alpha$ being close to 0, hence we refer to this notion as restricted weak agnostic learning. In the realizable setting where $\alpha=1$, we get $\beta=1$.
Intuition for why this is the right bound on $\beta$ can again be obtained from the geometric setting, where $q^{*},q,c$ are vectors in $\{\pm 1\}^{n}$. The condition $\langle q,c\rangle\geq\alpha$ translates to $d_{H}(q,c)\leq(1-\alpha)/2$, and similarly for $d_{H}(q^{*},c)$. By the triangle inequality, we must have
@@
While translating this intuition from the Boolean to the continuous setting is by no means immediate, we show that 2p12p-1 is indeed a $\beta=2\alpha-1$-weak learner. The key step is to show the following upper bound on the correlation obtainable using any hypothesis in $\mathcal{C}$: if $p$ is $(\mathcal{C},\tau)$-multiaccurate, then
@@
Thus a calibrated predictor that deviates noticeably from random guessing (saying $1/2$ all the time) gives rise to an agnostic learner. But if we add the assumption that $p$ is $\mathcal{C}$-multiaccurate, and that there exists $c\in\mathcal{C}$ so that $\mathrm{cor}(\mathbf{y},c(\mathbf{x}))\geq\alpha$, then $p$ can no longer predict $1/2$ all the time. This is proved in Lemma 4.6 which shows that for every $c\in\mathcal{C}$,
@@
We note that this result can also be shown using the loss OI technique of [17], where it is applied to show agnostic learning for Boolean function classes $\mathcal{C}$ (where correlation coincides with 0-11 loss). But we find the proof above more intuitive as it highlights how multiaccuracy and calibration complement each other.